### v1

# Predictive Inorganic Synthesis based on Machine Learning using Small Data sets: a case study of size-controlled Cu Nanoparticles

Abstract
Copper nanoparticles (Cu NPs) have a broad applicability, yet their synthesis is sensitive to subtle changes in reaction parameters. This sensitivity, combined with the time- and resource-intensive nature of experimental optimization, poses a major challenge in achieving reproducible and size-controlled synthesis. While Machine Learning (ML) shows promise in materials research, its application is often limited by scarcity of large high-quality experimental data sets. This study explores ML to predict the size of Cu NPs from microwave-assisted polyol synthesis using a small data set of 25 in-house performed syntheses. Latin Hypercube Sampling is used to efficiently cover the parameter space while creating the experimental data set. Ensemble regression models, built with the AMADEUS framework,[AMADEUS2020, vanpoucke:JournalApplPhys2020] successfully predict particle sizes with high accuracy (), outperforming classical statistical approaches (). Overall, this study highlights that, for lab-scale synthesis optimization, high-quality small datasets combined with classical, interpretable ML models outperform traditional statistical methods and are fully sufficient for quantitative synthesis prediction. This approach provides a sustainable and experimentally realistic pathway toward data-driven inorganic synthesis design.
1 Introduction
Metal nanoparticles (NPs) play a central role in a broad range of applications, including catalysis, electronics, and biomedicine, due to their unique size- and shape-dependent properties.[daniel:ChemInform2004, zhai:AdvancedFunctMat2010, gawande:ChemRev2016, pankhurst:JournalPhysD2003] However, controlling these properties requires very precise control of various NP characteristics, including their morphology (i.e., size and shape), composition, and surface chemistry down to the nanometer scale. Achieving this level of control remains one of the biggest challenges in NP synthesis, as small variations in the synthesis can lead to final particles with entirely different properties.[gawande:ChemRev2016, baig:MatAdv2021] Traditionally, synthesis optimization relies on iterative, empirical approaches that are slow, resource-intensive, and often lack mechanistic insight.[baig:MatAdv2021] To overcome these limitations, more systematic, data-driven strategies are needed, and recent advances in artificial intelligence (AI) and machine learning (ML) offer promising opportunities to accelerate the development of reproducible and tailored NP syntheses.[baig:MatAdv2021, baum:JournalChemInfMod2021, butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023, agrawal:APLMaterials2016]
By uncovering hidden correlations in complex parameter spaces, ML can guide experimental design and reduce the number of required experiments.[butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023] In materials science, ML is now routinely used to uncover the structure–property relationships of materials, thereby accelerating the discovery of new materials with specific properties.[butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023] However, for synthesis optimization, the key challenge lies in predicting how synthesis parameters determine the resulting crystal structure or particle morphology. These synthesis–structure relationships are far less explored, yet crucial for rational synthesis design. As a result, recent data-driven approaches are shifting towards models that learn synthesis–structure relationships to accelerate synthesis optimization.[agrawal:APLMaterials2016, vanpoucke:PolymerInt2022, williamson:InorgChem2023] Unlike traditional one-variable-at-a-time (OVAT) approaches, which are slow and fail to capture the multidimensional correlations that govern synthesis outcomes, ML can account for such interactions and thereby accelerate synthesis optimization.[williamson:InorgChem2023, lah:EngAnalysisBoundElements2024] However, while large data sets underpin many ML successes [Huo2022MLsynthesis, merchant-2023, curtarolo-2013], they are rarely feasible in laboratory-scale chemistry, where the creation of each data point requires a significant amount of time and resources. In the specific case of inorganic NP synthesis, literature-based data sets are often inconsistent and biased, as failed experiments remain unpublished[chen:AdvEnergyMat2020, raccuglia:Nature2016, ma:NPJCompMat2024], as shown by Huo et al.[Huo2022MLsynthesis], who expected up to % of the values to be incorrect. As such, robust ML workflows for small, high-quality data sets often remain underexplored.
ML can identify which synthesis parameters most strongly influence inorganic NP outcomes, as demonstrated by Guda et al.[guda:JournalPhysChemC2023], who trained ML models on 30 data points to classify the results of gold NP synthesis (success, shape, and formation kinetics) based on the precursor concentrations, reducing agent, and surfactant. ML also enables closed-loop experimentation, where ML models suggest subsequent experiments based on previous results, accelerating optimization and supporting high-throughput synthesis.[tao:AdvFunctMat2021, epps:AdvMat2020] Pellegrino et al.[pellegrino:ScientificRep2020] trained neural networks (NNs), with 20 initial data points, to predict the size, dispersion, and aspect ratio of NP, identifying optimal synthesis parameters through reverse engineering. Similarly, Williams et al.[williams:AppliedNano2023] combined Design of Experiments (DoE) and random forests to accurately predict the yield of Ag nanowires in a polyol-based flow reactor, based on 31 samples.
In the current work, the prediction of copper (Cu) NP size as function of synthesis parameters is explored. Cu NPs hold significant promise for various catalytic processes, such as electro-catalysis and photo-catalysis, due to their abundance, cost-effectiveness, and versatile properties.[guo:AngewandteChemIE2014, reske:JACS2014, gawande:ChemRev2016] For example, Reske et al.[reske:JACS2014] showed that decreasing the size of Cu NPs significantly improves their catalytic activity. Different synthesis methods give rise to different types of Cu NPs, which may differ in particle size, morphology, and extent of oxidation. While metallic Cu NPs are nominally (), they often undergo (partial) surface oxidation depending on the synthesis conditions, leading to () phases of varying composition.[gawande:ChemRev2016] The wide variety of possible synthesis outcomes, as well as the large parameter space, complicates the development of robust synthesis procedures to obtain specific properties.
Though ML is a tool with great potential for Cu NP synthesis, two challenges need to be tackled. The first challenge is finding a good quality data set to train the ML model, as a model can only be as good as its training data.[chen:AdvEnergyMat2020, raccuglia:Nature2016, ma:NPJCompMat2024] To tackle this challenge, a consistent and qualitative data set is built in-house. The second challenge is overcoming the data hungry nature of typical ML methods. Since the goal is to make synthesis optimization more efficient using ML, it makes no sense to first conduct hundreds of experiments that can serve as data to train the model. Therefore, a small data set of only a few experiments is used.[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022] Finally, within the context of lab-scale experiments, we emphasize interpretable ML: the model should not only predict outcomes, but also provide insight into the role of different synthesis parameters.
In this work, we develop an interpretable ML model for Cu NP synthesis based on a small, high-quality data set, aiming to accelerate synthesis optimization while providing fundamental insights into parameter effects. In addition, performance of the interpretable ML model is also compared with the more classical DoE approach.
2 Methodology
Within this work, we follow a four step workflow (Figure 1), of which the latter three can form a closed-loop experimentation methodology. The initial set of parameters for the data set is selected using Latin Hypercube Sampling (LHS) to achieve optimal parameter space coverage.[borisut:Processes2023] The sampled syntheses are performed in the lab via MW-assisted polyol synthesis. The resulting dispersions are analyzed via dynamic light scattering (DLS) to determine the particle size of the Cu NPs, while the surface plasmon resonance (SPR) peak is determined using UV-Vis measurements. The effective synthesis parameters111These are the true experimental values corresponding to the LHS. and corresponding particle sizes are combined into the data set for the ML modeling. In a closed-loop approach, the results of the ML models can be used to assist in further experimental syntheses, or further validation of the ML model.
2.1 Experimental synthesis
2.1.1 Materials
For all experimental syntheses, absolute ethanol (% analytical grade) obtained from Fisher Chemical and copper(II) acetate monohydrate (, ACS reagent), tetramethylammonium hydroxide (TMAH, wt. % in ) and ethylene glycol (%) obtained from Sigma Aldrich are used. All reagents are used as received without further purification.
2.1.2 Sampling synthesis parameters
LHS is used to obtain an optimized data set with minimal correlation between the various experiments while providing an optimal coverage of the parameter space under investigation.[teichert:DaltonTrans2018, borisut:Processes2023, braham:DaltonTrans2020] In this work, the parameter space comprises three dimensions:
1.
Cu precursor concentration (1 mM to 10 mM)
2.
Reaction time (1 minute to 20 minutes)
3.
Reaction temperature (175 ∘C to 200 ∘C)
Using LHS, 25 experiments are selected that fill this parameter space as optimally as possible.
2.1.3 Cu NPs synthesis
The Cu NPs are synthesized via the microwave (MW)-assisted polyol route[fievet:ChemSocRev2018, blosi:JournalNPResearch2010, teichert:DaltonTrans2018, carroll:JournalPhysChemC2011, torras:CrystalGrowth2021] in a CEM Discover SP MW reactor. First, the amount of is weighed into the reaction vial. It is then dissolved in 5 ml of ethylene glycol under sonication. After the Cu precursor has completely dissolved, the tetramethylammonium hydroxide (TMAH) is added to the solution and further mixed and degassed under sonication. TMAH is consistently used at a molar ratio of relative to the Cu precursor. The reaction vessel is then placed in the microwave reactor, with the head space filled with inert argon up to a pressure of 5 bar, and is first stirred for minutes at to ensure complete dissolution of the precursor. Afterwards, the temperature is raised to the desired reaction temperature with continuous stirring and maintained for the desired reaction time. After the reaction time has passed, the reaction is quenched by actively cooling the vessel with compressed air. A brown-red suspension is obtained, which is typical for Cu NPs.[blosi:JournalNPResearch2010] To collect the Cu NPs, the reaction mixture is diluted with ml of ethanol and centrifuged at rpm for minutes. Subsequently, the supernatant is removed, and the resulting NPs are again dispersed in ml of ethanol.
2.1.4 Characterization
The SPR of the synthesized Cu NPs in the ethylene glycol dispersion is determined by UV-Vis measurements with the Agilent Cary 5000 spectrophotometer equipped with a dual-beam configuration. Measurements are performed in a quartz cuvette ( mm path length) over a wavelength range of nm. Baseline correction is applied using the ethylene glycol solvent as reference.
The size of the synthesized Cu NPs is determined as the hydrodynamic diameter via DLS measurements with a Zetasizer Nano ZS, equipped with a nm He–Ne laser. For these measurements, the ethanolic Cu dispersions are sonicated for at least 10 minutes. Subsequently, five drops of the NP dispersion are diluted with ethanol to a total volume of approximately mL, ensuring the disposable polystyrene cuvette is filled above the required optical path.
2.2 Computational modeling
Two modeling tasks are performed on the data set: a regression task based on ensemble regression ML, implemented in the AMADEUS framework,[AMADEUS2020, vanpoucke:JournalApplPhys2020] and these results are compared with classical DoE[shahriari:ProceedingsIEEE2015], to gain insight in the potential of ML.
2.2.1 Ensemble regression models
Since ML using small data sets is very sensitive to the selected training data, and hyperparameter tuning suffers from this volatility[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022], we make use of the scikit-learn-based AMADEUS framework which was designed to deal with small data sets (Figure 2).[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022, pedregosafabianJournalMLResearch2011, AMADEUS2020] Pasting-type ensemble models consisting of base instances are created, using either linear or polynomial base models in combination with LASSO regularization.[tibshirani:RoyalStatSocB1996] Each of the instances is trained on a different randomly selected 80/20 split (in-bag/out-of-bag (OOB)) of the complete data set. The resulting final ensemble model is a single base instance equal to the average of the 100 base model instances.
To determine and compare the accuracy and generalization of the various models, root-mean-squared error (RMSE), mean absolute error (MAE) and coefficient of determination () are used as performance metrics. These errors can be determined on the in-bag and OOB data, with the difference between the two providing an indicator on model overfitting, while the individual values of the OOB error are quality indicators of the models’ generalization. A low training error and high OOB error are very strong indicators of a model that is overfitting. The average model quality (MAE, RMSE and ) is determined by having the final ensemble model make predictions for all data points in the entire data set, and correlates very well with the OOB quality measure.
We first examine the impact of multi-modal data in our dataset. For this purpose, three data sets are created:
1.
Data set A: All data points are included in the data set, and the predicted particle size is the smallest particle size present in the DLS particle-size distribution.
2.
Data set B: All data points are included in the data set, and the predicted particle size corresponds to the peak with the highest percentage contribution in the DLS particle-size distribution.
3.
Data set C: Only the data points with a mono-modal DLS particle-size distribution are included in the data set.
For each of these data sets, ensembles of polynomial regression models of order to with LASSO regularization () are trained. The performance metrics of the best performing ensemble model are presented in Table 1, and show that data set C performs significantly better. Therefore, only data set C will be considered further. The trained models in this study thus focus on predicting the particle size for a mono-modal distribution (cf. SI Figures 9, 10 and 11) and therefore cannot predict situations where a multi-modal distribution is obtained.
| polynomial degree | RMSE | MAE | ||
| Data set A | 3 | |||
| Data set B | 3 | |||
| Data set C | 3 |
When using a small data set, the growing number of features in polynomial models quickly leads to overfitting. Because the hyperparameter tuning of the regularization strength (LASSO regularization) presented itself to be too volatile for automatic optimization, it is selected based on visual comparison of four values ( and ). Ensemble models with polynomials up to degree eight as base model are trained. Analysis of the ensemble model MAE shows an overall optimal performance for (cf. SI Figure 5). Furthermore, the hyperparameter behavior shows no further improvement above polynomial degree three. Therefore, no features above degree three are considered, further reducing overfitting.
A further measure employed to reduce overfitting, is consideration of what one could call ensemble importance.[vanpoucke:PolymerInt2022, VerdingP:AdvMaterTech2025] The AMADEUS framework tracks how often a polynomial feature is retained (i.e. has non-zero coefficient) within the ensemble, after regularization of the features in each of the model instances. This provides a measure of importance of the features with regard to a well-generalizing model. A feature with an ensemble importance of % has a non-zero coefficient in all model instances in the ensemble. If a feature has an ensemble importance of %, this means that in % of the model instances in the ensemble, the coefficient for this feature is zero. Based on the trends in ensemble importance, seven features are selected (cf. SI Figure 6):
| 1. 2. T 3. 4. | 5. 6. 7. |
Since , T, and time are the base parameters used during synthesis, they are all retained as features. and are included as they play relatively important roles in both the second and third degree models. The interaction terms and are included, given their high importance in the second-degree polynomial; higher-order interaction terms are not included to limit model complexity and the probability of overfitting. Using these features, a linear regression model is trained, and iteratively, features are further engineered to obtain the optimal feature set.
2.2.2 Classical statistical methods (DoE)
Using DoE[shahriari:ProceedingsIEEE2015] software JMP Pro (SAS Institute Inc.), a linear regression model is fitted for data set C, including terms up to the second degree. Insignificant terms are removed until only significant terms remain. In contrast to machine learning approaches, the available data set is not split into training and test data; all data points are used to fit the regression model. The performance of the model is determined by calculating the expected size for all data points using the equation obtained, and thus should be compared to quality estimates on the in-bag data set.
3 Results and Discussion
| Training | OOB | Average | ||||||
| # Features | RMSE [nm] | MAE [nm] | RMSE [nm] | MAE [nm] | RMSE [nm] | MAE [nm] | ||
| 1st generation | 19 | 22.17 | 16.05 | 86.52 | 68.51 | 32.64 | 22.53 | 0.75 |
| 2nd generation | 7 | 29.40 | 20.47 | 51.85 | 41.36 | 32.61 | 22.28 | 0.75 |
| 3rd generation | 11 | 27.78 | 19.79 | 63.66 | 52.18 | 33.07 | 23.35 | 0.74 |
| 4th generation | 6 | 30.57 | 22.39 | 49.90 | 40.92 | 33.03 | 23.81 | 0.74 |
| DoE | 2 | - | - | - | - | 40.91 | 33.54 | 0.60 |
3.1 Cu NP syntheses
The DLS measurements show a peak above nm for five samples. However, these peaks are not included in the analysis. Due to their very low intensity (), it can be assumed that these results are due to contamination, such as dust particles, during the measurements. The synthesis yields Cu NPs with a particle size between nm and nm. The DLS shows a bi-modal and tri-modal particle-size distribution for five and two samples, respectively. The remaining samples show a mono-modal particle-size distribution. Analysis of the DLS intensity-size plots allows us to identify the smallest particle size present. The median of these particle sizes is nm over all samples.
Based on the SPR location, the effective formation of Cu metal () NPs is verified for the different samples. The different UV-Vis spectra show a clear SPR peak between nm and nm for most samples, which is a strong indication of the presence of NPs.[blosi:JournalNPResearch2010] Samples , , and do not show a clear SPR feature, which may be due to low particle concentration, oxidation[vargas-urbano:Materials2022], or the presence of ultra small particles (), which do not exhibit a well-defined SPR; for the other samples, at least some is present. Given the limited yield of the syntheses, no X-ray diffraction measurements could be performed to provide a conclusive answer, but the combination of the red-brown color of the solution and the presence of a clear SPR peak in the UV-Vis spectra strongly indicate that NPs have indeed formed. To validate this conclusion, Selected Area Electron Diffraction analysis is conducted on sample 22 using a transmission electron microscope operated at an accelerating voltage of 120 kV. The resulting diffraction rings are indexed to the face-centered cubic structure of copper, with prominent rings corresponding to the (111), (200), and (220) planes. The primary diffraction rings are assigned the following spacings: Å for the (111) plane, Å for the (200) plane, and Å for the (220) plane. These values agree with the standard values for FCC copper (JCPDS card no. 04-0836)[icdd2001], confirming the formation of NPs.
3.2 Quantitative size prediction
3.2.1 Ensemble regression
Ensemble regression models are created to predict Cu NP size based on the synthesis parameters. Different generations of models are trained to investigate the influence of different features on model performance.
As first generation ensemble model, we consider the complete third-degree polynomial model with LASSO regularization (). This model has an of , and RMSE and MAE of nm and nm, respectively. Despite the very good performance of this first generation model, its main disadvantage is the use of an extensive set of features. With our data set consisting of only 18 samples, the risk of overfitting is significant. In addition, a larger feature set results in a more complex model and, therefore, higher computational cost for inference. With this in mind, the following generations are aimed at reducing the model size drastically to improve generalization and reduce computational cost.
In the second generation, the complexity of the model is strongly reduced by using only the seven features (, , T, , , , ) selected based on the ensemble importance analysis. The ensemble average shows no significant loss in performance compared to the first generation (cf., Table 2), despite the significant reduction in the number of features used. In addition, the training and OOB errors of the ensemble show that the second generation is more generalizable as their difference becomes smaller. The training error of the second generation increases slightly (MAE of nm compared to nm), while the OOB error decreases drastically (MAE goes from nm to nm). The smaller difference between training and OOB error indicates that the model is overfitting less and generalizes better.
In the third generation model, the feature set is expanded with four additional features, aimed at the investigation of the influence of non-polynomial features (based on domain knowledge) on the model performance.
1.
is added because often has a high and consistent ensemble importance in the investigation of different polynomial models where it occurs in higher powers.
2.
can be interpreted as a measure of the amount of cations available per unit of time during the growth of the particles.
3.
provides a measure of the total exposure of the reduction medium to Cu ions.
4.
as a temperature-normalized reaction time. Although this parameter does not represent the classical thermal budget (), it provides a simple heuristic descriptor that captures part of the interplay between time and temperature in a single scalar.
This more complex feature set, gives rise to a slight decline in the performance of the model due to stronger overfitting, leading to poorer generalization. Given the small difference in performance between this and the previous generation, a comprehensive analysis of the ensemble and dominance importance[Azen:PsyMeth2003, Ceder:ChemMater2022] of the features is carried out. Based on this analysis features are selected, aimed at a good balance between sufficient complexity for good predictive quality and a reduced size while retaining good generalization, which resulted in our fourth generation model.
The dominance importance and ensemble importance are considered in an iterative fashion, as done in previous work of one of the authors.[vanpoucke:PolymerInt2022, VerdingP:AdvMaterTech2025] An exhaustive set of sub-models are trained, with one, two or three features (from the full feature list of the third generation model) systematically disabled. This allows us to determine the influence of each individual feature on the model’s performance. Five features are found to have a low importance (%) on average over all models. Therefore, only the six features with an average ensemble importance % are kept for the fourth generation (cf. SI Table IX).
| 1. T 2. 3. | 4. 5. 6. |
The predicted particle sizes resulting from our fourth generation ensemble regression model are shown in figure 3. In this model, each of the six features has an ensemble importance of % (LASSO regularization, ). The quality of the model is only very slightly below the 19-feature model, showing a significant improvement with respect to overfitting (cf. Table 2). Because the size of the model has been greatly reduced, the MAE for the OOB predictions decreases from nm for the third generation to nm for the fourth generation, a marked improvement. Since the average of the ensemble model is equivalent to a single model instance with averaged coefficients [vanpoucke:JournalApplPhys2020], our fourth generation model can be formulated as follows:
In Equation 1, all feature values are standardized prior to training, meaning the coefficients apply to standardized inputs (cf. SI Table X). When standardized experimental feature values are inserted into Equation 1, the model directly returns the predicted particle size in nm.
Equation 2 gives the non-standardized form of the fourth generation model. Here, in molar, in minutes and T in can be used directly to obtain the particle size in nm.
3.2.2 Statistical approach (DoE)
In the classical statistical approach, only the concentration of the Cu precursor appears to significantly influence the particle size, resulting in a quadratic model shown in figure 4 and given by Equation 3. Here, [Cu] refers to the non-standardized precursor concentration in molar units, and the equation returns the Cu NP size directly in nm. As seen in Table 2 the quality is significantly lower than the ML regression models. In addition, it is unclear whether the model will generalize successfully, as it is based on the entire data set.
| Cu NP size | ||||
Since the ensemble regression models are built on random subsets of the entire data set, it can be expected that they will generalize better than the classical regression model that was trained on the entire data set. Because individual data points have a large impact in small datasets, models trained on the full dataset are highly sensitive to noise and outliers. Training multiple model instances on random subsets and averaging their predictions reduces this sensitivity, yielding more accurate and robust models.[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022] On the other hand, the classical regression model consists of only two terms containing the precursor concentration, compared to the six features in the ensemble regression model, which results in a simpler model and therefore lower computational cost. Comparison of the ML and DoE models indicates that Cu concentration is an important factor, as both contain it as a polynomial term. In addition, the influence of time and temperature are of relevance as well (as known from experiment) which is captured by the ML models only.
Since the DoE model only retained two features, we attempted to further reduce the fourth generation model by removal of additional features, based on dominance importance analyses (cf. SI Table XI). First was removed, resulting in a slight performance reduction. dropped from to and the average MAE increased from nm to nm. On the OOB data, both models have almost the same MAE, so despite the slightly lower performance there is no increase in generalization. Next, also was removed, as it was assumed to be covered by the feature. Now the overall performance further dropped with an of and an average MAE of nm. In this case, generalization does increase significantly with a training MAE of nm and an OOB MAE of nm. Finally, also was removed from the feature list, to arrive at a three feature model. Again, the generalization of the model increases with a training MAE of nm and an OOB MAE of nm, while maintaining similar average model performance with an of and an average MAE of nm.
This demonstrates the delicate balance between average model performance and overall model generalization. Reducing model size does give significant improvements in generalization, but comes with a cost in average performance. While the fourth generation model with six features had an average MAE of only nm, this increased to nm for the three feature model, getting closer to the nm MAE of the two feature DoE model.
3.2.3 Synthesis prediction: model validation
| Sample | Targeted [nm] | Obtained [nm] | DoE prediction [nm] |
| 26 | ) | 162.67 | |
| 27 | ) | 153.30 | |
| 28 | ) | 220.86 | |
| 29 | ) | 225.19 | |
| 30 | ) | 251.43 | |
| 31 | ) | 250.39 |
For three different target sizes ( nm, nm and nm) Equation 2 is used to suggest two possible syntheses routes for each (cf. SI Table XII). The suggested syntheses are performed using the same experimental procedure used for creating the original data set samples. For two of the targeted particle sizes one of the suggested syntheses routes closely matched the target size within experimental uncertainty (cf. samples & in Table 3). For the largest particles (samples & ) neither suggested set of parameters gave the expected results. It is interesting to note that for the incorrect results, the particles are always too small. The suggested synthesis are also used to predict particle size using the DoE model (Equation 3). This shows that the DoE model exhibits a flatter, less pronounced curvature in the explored parameter space compared to the ML model.
4 Conclusion
Despite the small size of the data set, it is possible to predict the particle size of Cu NPs quantitatively based on synthesis parameters using classical ensemble regression ML, achieving an of and an MAE of nm. For most data points, the predicted particle size falls within the experimental standard deviation. We were able to reduce the number of features in our final model to six, which improved generalization and reduced model size.
By in-house constructing a consistent data set, we overcome limitations associated with biased or incomplete literature data. The use of interpretable models allowed us to identify the synthesis parameters with the strongest influence on particle properties, providing fundamental insights alongside predictive power. As validation, the final model is used to suggest synthesis routes for specific target particle sizes. For two of the three target sizes, the suggested syntheses closely matched the experimental results, demonstrating that ML can effectively guide experimental synthesis even with a minimal data set.
Although this work focused on Cu NPs as a model system, the approach is broadly applicable to other materials. Future work could integrate such ML workflows into closed-loop experimentation and explore larger, more diverse data sets. Overall, this strategy contributes to more efficient, interpretable, and sustainable materials development.
Acknowledgment
The computational resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government–department WEWIS.
Supplemental Information
| Sample | Theoretical Cu conc.a [mM] | Effective Cu prec.b [mM] | TMAHc [l] | [min] | Temp.e [∘C] |
| 1 | 5.00 | 5.06 | 36.0 | 8 | 196 |
| 2 | 4.00 | 4.21 | 29.0 | 4 | 188 |
| 3 | 3.00 | 2.91 | 21.8 | 6 | 198 |
| 4 | 2.00 | 2.00 | 14.6 | 7 | 184 |
| 5 | 9.00 | 8.82 | 65.4 | 4 | 180 |
| 6 | 3.00 | 3.11 | 21.8 | 1 | 194 |
| 7 | 10.00 | 10.22 | 72.8 | 9 | 188 |
| 8 | 9.00 | 8.82 | 65.4 | 16 | 182 |
| 9 | 3.00 | 2.80 | 21.8 | 17 | 182 |
| 10 | 7.00 | 6.91 | 51.0 | 20 | 194 |
| 11 | 7.00 | 6.81 | 51.0 | 14 | 186 |
| 12 | 6.00 | 6.01 | 43.6 | 10 | 192 |
| 13 | 8.00 | 8.31 | 58.2 | 12 | 184 |
| 14 | 1.00 | 1.10 | 7.2 | 15 | 198 |
| 15 | 2.00 | 2.10 | 14.6 | 10 | 196 |
| 16 | 4.00 | 3.91 | 29.0 | 18 | 200 |
| 17 | 4.00 | 3.91 | 29.0 | 14 | 192 |
| 18 | 8.00 | 8.31 | 58.2 | 5 | 180 |
| 19 | 2.00 | 2.10 | 14.6 | 17 | 176 |
| 20 | 8.00 | 7.91 | 58.2 | 11 | 190 |
| 21 | 6.00 | 5.91 | 43.6 | 7 | 178 |
| 22 | 7.00 | 6.91 | 51.0 | 3 | 176 |
| 23 | 9.00 | 9.02 | 65.4 | 13 | 178 |
| 24 | 6.00 | 6.11 | 43.6 | 19 | 186 |
| 25 | 5.00 | 5.01 | 36.0 | 2 | 190 |
| Sample | Effective Cu conc.a [mM] | TMAHb [l] | [min] | Temp.d [∘C] | Sizee [nm] |
| 1 | 5.06 | 36.0 | 8 | 196 | 252.50 (59.37) |
| 2 | 4.21 | 29.0 | 4 | 188 | 228.40 (41.47) |
| 3 | 2.91 | 21.8 | 6 | 198 | 148.80 (25.56) |
| 4 | 2.00 | 14.6 | 7 | 184 | 99.90 (14.07) |
| 5 | 8.82 | 65.4 | 4 | 180 | 183.30 (25.42) |
| 6 | 3.11 | 21.8 | 1 | 194 | 160.70 (27.62) |
| 7 | 10.22 | 72.8 | 9 | 188 | 64.65 (8.80) |
| 8 | 8.82 | 65.4 | 16 | 182 | 40.85 (5.20) |
| 9 | 2.80 | 21.8 | 17 | 182 | 192.90 (28.88) |
| 10 | 6.91 | 51.0 | 20 | 194 | 237.90 (30.29) |
| 11 | 6.81 | 51.0 | 14 | 186 | 167.00 (31.15) |
| 12 | 6.01 | 43.6 | 10 | 192 | 60.10 (6.88) |
| 13 | 8.31 | 58.2 | 12 | 184 | 66.02 (11.94) |
| 14 | 1.10 | 7.2 | 15 | 198 | 102.90 (13.51) |
| 15 | 2.10 | 14.6 | 10 | 196 | 114.40 (17.64) |
| 16 | 3.91 | 29.0 | 18 | 200 | 157.90 (22.70) |
| 17 | 3.91 | 29.0 | 14 | 192 | 304.50 (52.21) |
| 18 | 8.31 | 58.2 | 5 | 180 | 59.09 (6.13) |
| 19 | 2.10 | 14.6 | 17 | 176 | 188.10 (26.04) |
| 20 | 7.91 | 58.2 | 11 | 190 | 50.36 (8.16) |
| 21 | 5.91 | 43.6 | 7 | 178 | 312.90 (55.82) |
| 22 | 6.91 | 51.0 | 3 | 176 | 271.60 (45.13) |
| 23 | 9.02 | 65.4 | 13 | 178 | 127.60 (18.55) |
| 24 | 6.11 | 43.6 | 19 | 186 | 58.69 (10.61) |
| 25 | 5.01 | 36.0 | 2 | 190 | 257.60 (48.85) |
| Sample | Effective Cu conc.a [mM] | TMAHb [l] | [min] | Temp.d [∘C] | Sizee [nm] |
| 1 | 5.06 | 36.0 | 8 | 196 | 252.50 (59.37) |
| 2 | 4.21 | 29.0 | 4 | 188 | 228.40 (41.47) |
| 3 | 2.91 | 21.8 | 6 | 198 | 148.80 (25.56) |
| 4 | 2.00 | 14.6 | 7 | 184 | 99.90 (14.07) |
| 5 | 8.82 | 65.4 | 4 | 180 | 183.30 (25.42) |
| 6 | 3.11 | 21.8 | 1 | 194 | 160.70 (27.62) |
| 7 | 10.22 | 72.8 | 9 | 188 | 204.00 (34.54) |
| 8 | 8.82 | 65.4 | 16 | 182 | 222.30 (31.62) |
| 9 | 2.80 | 21.8 | 17 | 182 | 192.90 (28.88) |
| 10 | 6.91 | 51.0 | 20 | 194 | 237.90 (30.29) |
| 11 | 6.81 | 51.0 | 14 | 186 | 167.00 (31.15) |
| 12 | 6.01 | 43.6 | 10 | 192 | 269.50 (60.33) |
| 13 | 8.31 | 58.2 | 12 | 184 | 66.02 (11.94) |
| 14 | 1.10 | 7.2 | 15 | 198 | 102.90 (13.51) |
| 15 | 2.10 | 14.6 | 10 | 196 | 114.40 (17.64) |
| 16 | 3.91 | 29.0 | 18 | 200 | 157.90 (22.70) |
| 17 | 3.91 | 29.0 | 14 | 192 | 304.50 (52.21) |
| 18 | 8.31 | 58.2 | 5 | 180 | 184.90 (26.97) |
| 19 | 2.10 | 14.6 | 17 | 176 | 188.10 (26.04) |
| 20 | 7.91 | 58.2 | 11 | 190 | 50.36 (8.16) |
| 21 | 5.91 | 43.6 | 7 | 178 | 312.90 (55.82) |
| 22 | 6.91 | 51.0 | 3 | 176 | 271.60 (45.13) |
| 23 | 9.02 | 65.4 | 13 | 178 | 127.60 (18.55) |
| 24 | 6.11 | 43.6 | 19 | 186 | 58.69 (10.61) |
| 25 | 5.01 | 36.0 | 2 | 190 | 257.60 (48.85) |
| Sample | Effective Cu conc.a [mM] | TMAHb [l] | [min] | Temp.d [∘C] | Sizee [nm] |
| 1 | 5.06 | 36.0 | 8 | 196 | 252.50 (59.37) |
| 2 | 4.21 | 29.0 | 4 | 188 | 228.40 (41.47) |
| 3 | 2.91 | 21.8 | 6 | 198 | 148.80 (25.56) |
| 4 | 2.00 | 14.6 | 7 | 184 | 99.90 (14.07) |
| 5 | 8.82 | 65.4 | 4 | 180 | 183.30 (25.42) |
| 6 | 3.11 | 21.8 | 1 | 194 | 160.70 (27.62) |
| 9 | 2.80 | 21.8 | 17 | 182 | 192.90 (28.88) |
| 10 | 6.91 | 51.0 | 20 | 194 | 237.90 (30.29) |
| 11 | 6.81 | 51.0 | 14 | 186 | 167.00 (31.15) |
| 14 | 1.10 | 7.2 | 15 | 198 | 102.90 (13.51) |
| 15 | 2.10 | 14.6 | 10 | 196 | 114.40 (17.64) |
| 16 | 3.91 | 29.0 | 18 | 200 | 157.90 (22.70) |
| 17 | 3.91 | 29.0 | 14 | 192 | 304.50 (52.21) |
| 19 | 2.10 | 14.6 | 17 | 176 | 188.10 (26.04) |
| 21 | 5.91 | 43.6 | 7 | 178 | 312.90 (55.82) |
| 22 | 6.91 | 51.0 | 3 | 176 | 271.60 (45.13) |
| 23 | 9.02 | 65.4 | 13 | 178 | 127.60 (18.55) |
| 25 | 5.01 | 36.0 | 2 | 190 | 257.60 (48.85) |
| Sample | Size_1a [nm] | Percentage_1b [%] | Size_2a [nm] | Percentage_2b [%] | Size_3a [nm] | Percentage_3b [%] | SPRd [nm] |
| 1 | 252.50 (59.37) | 99.9 | 5414.00 (647.80)c | 0.1 | - | - | 613 |
| 2 | 228.40 (41.47) | 99.9 | 5135.00 (775.60)c | 0.1 | - | - | 625 |
| 3 | 148.80 (25.56) | 100 | - | - | - | - | 612 |
| 4 | 99.90 (14.07) | 100 | - | - | - | - | 699 |
| 5 | 183.30 (25.42) | 100 | - | - | - | - | 597 |
| 6 | 160.70 (27.62) | 100 | - | - | - | - | 699 |
| 7 | 64.65 (8.80) | 29.1 | 204.00 (34.54) | 70.9 | - | - | 592 |
| 8 | 40.85 (5.20) | 33.1 | 222.30 (31.62) | 66.9 | - | - | 590 |
| 9 | 192.90 (28.88) | 100 | - | - | - | - | 609 |
| 10 | 237.90 (30.29) | 100 | - | - | - | - | 591 |
| 11 | 167.00 (31.15) | 98.9 | 5164.00 (744.50)c | 1.1 | - | - | 636 |
| 12 | 60.10 (6.88) | 27.1 | 123.70 (22.78) | 14.1 | 269.50 (60.33) | 58.8 | 622 |
| 13 | 66.02 (11.94) | 53.0 | 126.20 (18.36) | 4.7 | 259.90 (45.83) | 42.3 | 611 |
| 14 | 102.90 (13.51) | 100 | - | - | - | - | 618 |
| 15 | 114.40 (17.64) | 100 | - | - | - | - | 619 |
| 16 | 157.90 (22.70) | 100 | - | - | - | - | 593 |
| 17 | 304.50 (52.21) | 100 | - | - | - | - | 616 |
| 18 | 59.09 (6.13) | 30.4 | 184.90 (26.97) | 69.6 | - | - | 605 |
| 19 | 188.10 (26.04) | 99.7 | 5532.00 (608.70)c | 0.3 | - | - | 626 |
| 20 | 50.36 (8.16) | 63.5 | 233.10 (37.46) | 36.5 | - | - | 595 |
| 21 | 312.90 (55.82) | 100 | - | - | - | - | 609 |
| 22 | 271.60 (45.13) | 99.9 | 5380.00 (654.70)c | 0.1 | - | - | 601 |
| 23 | 127.60 (18.55) | 100 | - | - | - | - | 595 |
| 24 | 58.69 (10.61) | 56.4 | 247.00 (39.81) | 43.6 | - | - | 600 |
| 25 | 257.60 (48.85) | 100 | - | - | - | - | 597 |
| Feature | Average EI [%] | Minimum EI [%] | Maximum EI [%] |
| 84.5 | 10.0 | 100 | |
| T | 99.6 | 97.0 | 100 |
| time | 82.2 | 54.0 | 100 |
| 82.3 | 53.0 | 100 | |
| 80.0 | 43.0 | 100 | |
| 99.1 | 95.0 | 100 | |
| 98.0 | 85.0 | 100 | |
| 91.7 | 62.0 | 100 | |
| 95.1 | 84.0 | 100 | |
| 91.9 | 80.0 | 100 | |
| 84.7 | 63.0 | 100 |
| Feature | Mean | STDEV |
| T | ||
| T | MAE train | MAE OOB | RMSE train | RMSE OOB | |||||
| 1 | 1 | 1 | 1 | 0 | 1 | 24.45 | 37.69 | 32.47 | 45.57 |
| 1 | 0 | 1 | 1 | 0 | 1 | 26.21 | 36.86 | 34.73 | 44.36 |
| 1 | 0 | 1 | 1 | 0 | 0 | 26.66 | 34.43 | 35.33 | 42.58 |
| Sample | Cu conc.a [mM] | TMAHb [l] | [min] | Temp.d [∘C] | Targeted sizee [nm] | Obtained sizef [nm] |
| 26 | 8.82 | 64.0 | 18 | 175 | 115 | 70.78 (4.68) |
| 27 | 2.30 | 16.8 | 4 | 195 | 115 | 117.90 (32.06) |
| 28 | 7.51 | 54.6 | 16 | 175 | 220 | 235.60 (34.57) |
| 29 | 3.91 | 27.8 | 7 | 189 | 220 | 128.60 (49.45) |
| 30 | 5.51 | 40.0 | 15 | 175 | 275 | 165.10 (41.04) |
| 31 | 6.01 | 43.6 | 4 | 179 | 275 | 189.60 (59.50) |

### v2

@@
 Latin Hypercube Sampling is used to efficiently cover the parameter space while creating the experimental data set.
-Ensemble regression models, built with the AMADEUS framework,[AMADEUS2020, vanpoucke:JournalApplPhys2020] successfully predict particle sizes with high accuracy (), outperforming classical statistical approaches ().
-Overall, this study highlights that, for lab-scale synthesis optimization, high-quality small datasets combined with classical, interpretable ML models outperform traditional statistical methods and are fully sufficient for quantitative synthesis prediction.
-This approach provides a sustainable and experimentally realistic pathway toward data-driven inorganic synthesis design.
+Ensemble regression models successfully predict particle sizes with high accuracy (), outperforming classical statistical approaches ().
+Additionally, classification models using both random forests and Large Language Models (LLMs) are evaluated to distinguish between large and small particles.
+While random forests show moderate performance, LLMs offer no significant advantages under data-scarce conditions.
+Overall, this study demonstrates that carefully curated small data sets, paired with robust classical ML, can effectively predict the synthesis of Cu NPs and highlights that for lab-scale studies, complex models like LLMs may offer limited benefit over simpler techniques.
 1 Introduction
-Metal nanoparticles (NPs) play a central role in a broad range of applications, including catalysis, electronics, and biomedicine, due to their unique size- and shape-dependent properties.[daniel:ChemInform2004, zhai:AdvancedFunctMat2010, gawande:ChemRev2016, pankhurst:JournalPhysD2003] However, controlling these properties requires very precise control of various NP characteristics, including their morphology (i.e., size and shape), composition, and surface chemistry down to the nanometer scale.
-Achieving this level of control remains one of the biggest challenges in NP synthesis, as small variations in the synthesis can lead to final particles with entirely different properties.[gawande:ChemRev2016, baig:MatAdv2021] Traditionally, synthesis optimization relies on iterative, empirical approaches that are slow, resource-intensive, and often lack mechanistic insight.[baig:MatAdv2021] To overcome these limitations, more systematic, data-driven strategies are needed, and recent advances in artificial intelligence (AI) and machine learning (ML) offer promising opportunities to accelerate the development of reproducible and tailored NP syntheses.[baig:MatAdv2021, baum:JournalChemInfMod2021, butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023, agrawal:APLMaterials2016]
-By uncovering hidden correlations in complex parameter spaces, ML can guide experimental design and reduce the number of required experiments.[butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023] In materials science, ML is now routinely used to uncover the structure–property relationships of materials, thereby accelerating the discovery of new materials with specific properties.[butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023] However, for synthesis optimization, the key challenge lies in predicting how synthesis parameters determine the resulting crystal structure or particle morphology.
+Metal nanoparticles (NPs) play a central role in a broad range of applications, including catalysis, electronics, and biomedicine, due to their unique size- and shape-dependent properties.[14, 48, 18, 30] However, controlling these properties requires very precise control of various NP characteristics, including their morphology (i.e., size and shape), composition, and surface chemistry down to the nanometer scale.
+Achieving this level of control remains one of the biggest challenges in NP synthesis, as small variations in the synthesis can lead to final particles with entirely different properties.[18, 4] Traditionally, synthesis optimization relies on iterative, empirical approaches that are slow, resource-intensive, and often lack mechanistic insight.[4] To overcome these limitations, more systematic, data-driven strategies are an interesting route to explore.
+Indeed, recent advances in artificial intelligence (AI) and machine learning (ML) offer promising opportunities to accelerate the development of reproducible and tailored NP syntheses.[4, 5, 9, 35, 34, 21, 2]
+By uncovering hidden correlations in complex parameter spaces, ML can guide experimental design and reduce the number of required experiments.[9, 35, 34, 21] In materials science, ML is now routinely used to uncover the structure–property relationships of materials, thereby accelerating the discovery of new materials with specific properties.[9, 35, 34, 21] However, for synthesis optimization, the key challenge lies in predicting how synthesis parameters determine the resulting crystal structure or particle morphology.
 These synthesis–structure relationships are far less explored, yet crucial for rational synthesis design.
-As a result, recent data-driven approaches are shifting towards models that learn synthesis–structure relationships to accelerate synthesis optimization.[agrawal:APLMaterials2016, vanpoucke:PolymerInt2022, williamson:InorgChem2023] Unlike traditional one-variable-at-a-time (OVAT) approaches, which are slow and fail to capture the multidimensional correlations that govern synthesis outcomes, ML can account for such interactions and thereby accelerate synthesis optimization.[williamson:InorgChem2023, lah:EngAnalysisBoundElements2024] However, while large data sets underpin many ML successes [Huo2022MLsynthesis, merchant-2023, curtarolo-2013], they are rarely feasible in laboratory-scale chemistry, where the creation of each data point requires a significant amount of time and resources.
-In the specific case of inorganic NP synthesis, literature-based data sets are often inconsistent and biased, as failed experiments remain unpublished[chen:AdvEnergyMat2020, raccuglia:Nature2016, ma:NPJCompMat2024], as shown by Huo et al.[Huo2022MLsynthesis], who expected up to % of the values to be incorrect.
+As a result, recent data-driven approaches are shifting towards models that learn synthesis–structure relationships to accelerate synthesis optimization.[2, 42, 47] Unlike traditional one-variable-at-a-time (OVAT) approaches, which are slow and fail to capture the multidimensional correlations that govern synthesis outcomes, ML can account for such interactions and thereby accelerate synthesis optimization.[47, 26] However, while large data sets underpin many ML successes [23, 28, 13], they are rarely feasible in laboratory-scale chemistry, where the creation of each data point requires a significant amount of time and resources.
+In the specific case of inorganic NP synthesis, literature-based data sets are often inconsistent and biased, as failed experiments remain unpublished[11, 32, 27], as shown by Huo et al.[23], who expected up to % of the values to be incorrect.
 As such, robust ML workflows for small, high-quality data sets often remain underexplored.
-ML can identify which synthesis parameters most strongly influence inorganic NP outcomes, as demonstrated by Guda et al.[guda:JournalPhysChemC2023], who trained ML models on 30 data points to classify the results of gold NP synthesis (success, shape, and formation kinetics) based on the precursor concentrations, reducing agent, and surfactant.
-ML also enables closed-loop experimentation, where ML models suggest subsequent experiments based on previous results, accelerating optimization and supporting high-throughput synthesis.[tao:AdvFunctMat2021, epps:AdvMat2020] Pellegrino et al.[pellegrino:ScientificRep2020] trained neural networks (NNs), with 20 initial data points, to predict the size, dispersion, and aspect ratio of NP, identifying optimal synthesis parameters through reverse engineering.
-Similarly, Williams et al.[williams:AppliedNano2023] combined Design of Experiments (DoE) and random forests to accurately predict the yield of Ag nanowires in a polyol-based flow reactor, based on 31 samples.
+ML can identify which synthesis parameters most strongly influence inorganic NP outcomes, as demonstrated by Guda et al.[19], who trained ML models on 30 data points to classify the results of gold NP synthesis (success, shape, and formation kinetics) based on the precursor concentrations, reducing agent, and surfactant.
+ML also enables closed-loop experimentation, where ML models suggest subsequent experiments based on previous results, accelerating optimization and supporting high-throughput synthesis.[37, 15] Pellegrino et al.[31] trained neural networks (NNs), with 20 initial data points, to predict the size, dispersion, and aspect ratio of NP, identifying optimal synthesis parameters through reverse engineering.
+Similarly, Williams et al.[46] combined Design of Experiments (DoE) and random forests to accurately predict the yield of Ag nanowires in a polyol-based flow reactor, based on 31 samples.
 In the current work, the prediction of copper (Cu) NP size as function of synthesis parameters is explored.
-Cu NPs hold significant promise for various catalytic processes, such as electro-catalysis and photo-catalysis, due to their abundance, cost-effectiveness, and versatile properties.[guo:AngewandteChemIE2014, reske:JACS2014, gawande:ChemRev2016] For example, Reske et al.[reske:JACS2014] showed that decreasing the size of Cu NPs significantly improves their catalytic activity.
+Cu NPs hold significant promise for various catalytic processes, such as electro-catalysis and photo-catalysis, due to their abundance, cost-effectiveness, and versatile properties.[20, 33, 18] For example, Reske et al.[33] showed that decreasing the size of Cu NPs significantly improves their catalytic activity.
 Different synthesis methods give rise to different types of Cu NPs, which may differ in particle size, morphology, and extent of oxidation.
-While metallic Cu NPs are nominally (), they often undergo (partial) surface oxidation depending on the synthesis conditions, leading to () phases of varying composition.[gawande:ChemRev2016] The wide variety of possible synthesis outcomes, as well as the large parameter space, complicates the development of robust synthesis procedures to obtain specific properties.
+While metallic Cu NPs are nominally , they often undergo (partial) surface oxidation depending on the synthesis conditions, leading to phases of varying composition.[18] The wide variety of possible synthesis outcomes, as well as the large parameter space, complicates the development of robust synthesis procedures to obtain specific properties.
 Though ML is a tool with great potential for Cu NP synthesis, two challenges need to be tackled.
-The first challenge is finding a good quality data set to train the ML model, as a model can only be as good as its training data.[chen:AdvEnergyMat2020, raccuglia:Nature2016, ma:NPJCompMat2024] To tackle this challenge, a consistent and qualitative data set is built in-house.
+The first challenge is finding a good quality data set to train the ML model.[11, 32, 27] To tackle this challenge, a consistent and qualitative data set is built in-house.
 The second challenge is overcoming the data hungry nature of typical ML methods.
-Since the goal is to make synthesis optimization more efficient using ML, it makes no sense to first conduct hundreds of experiments that can serve as data to train the model.
-Therefore, a small data set of only a few experiments is used.[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022] Finally, within the context of lab-scale experiments, we emphasize interpretable ML: the model should not only predict outcomes, but also provide insight into the role of different synthesis parameters.
+Since the goal is to make synthesis optimization more efficient using ML, relying on hundreds of preliminary experiments to generate training data would be impractical in a lab-scale setting, both in terms of time and resource consumption.
+Instead, we focus on constructing a carefully selected, information-rich dataset that maximizes experimental variation while minimizing experimental effort.[29, 42] Finally, within the context of lab-scale experiments, we emphasize interpretable ML: the model should not only predict outcomes, but ideally also provide insight into the role of different synthesis parameters.
 In this work, we develop an interpretable ML model for Cu NP synthesis based on a small, high-quality data set, aiming to accelerate synthesis optimization while providing fundamental insights into parameter effects.
 In addition, performance of the interpretable ML model is also compared with the more classical DoE approach.
+Recent work by Zheng et al.[49], which applies large language models (LLMs) to synthesis protocols reported in the MOF literature, served as inspiration to examine the potential of LLMs for synthesis prediction.
+Motivated by this, we investigate whether LLMs can provide meaningful predictive performance in a small-data experimental setting and benchmark their performance against a classic ML approach.
 2 Methodology
 Within this work, we follow a four step workflow (Figure 1), of which the latter three can form a closed-loop experimentation methodology.
-The initial set of parameters for the data set is selected using Latin Hypercube Sampling (LHS) to achieve optimal parameter space coverage.[borisut:Processes2023] The sampled syntheses are performed in the lab via MW-assisted polyol synthesis.
+The initial set of parameters for the data set is selected using Latin Hypercube Sampling (LHS) to achieve optimal parameter space coverage.[7] The sampled syntheses are performed in the lab via MW-assisted polyol synthesis.
 The resulting dispersions are analyzed via dynamic light scattering (DLS) to determine the particle size of the Cu NPs, while the surface plasmon resonance (SPR) peak is determined using UV-Vis measurements.
-The effective synthesis parameters111These are the true experimental values corresponding to the LHS.
-and corresponding particle sizes are combined into the data set for the ML modeling.
+The effective synthesis parameters[1] and corresponding particle sizes are combined into the data set for the ML modeling.
 In a closed-loop approach, the results of the ML models can be used to assist in further experimental syntheses, or further validation of the ML model.
@@
 2.1.2 Sampling synthesis parameters
-LHS is used to obtain an optimized data set with minimal correlation between the various experiments while providing an optimal coverage of the parameter space under investigation.[teichert:DaltonTrans2018, borisut:Processes2023, braham:DaltonTrans2020] In this work, the parameter space comprises three dimensions:
+LHS is used to obtain an optimized data set with minimal correlation between the various experiments while providing an optimal coverage of the parameter space under investigation.[38, 7, 8] In this work, the parameter space comprises three dimensions:
 1.
@@
 Reaction temperature (175 ∘C to 200 ∘C)
-Using LHS, 25 experiments are selected that fill this parameter space as optimally as possible.
+Using LHS, 25 experiments are selected that fill this parameter space as optimally as possible. (cf. SI Table V)
 2.1.3 Cu NPs synthesis
-The Cu NPs are synthesized via the microwave (MW)-assisted polyol route[fievet:ChemSocRev2018, blosi:JournalNPResearch2010, teichert:DaltonTrans2018, carroll:JournalPhysChemC2011, torras:CrystalGrowth2021] in a CEM Discover SP MW reactor.
+The Cu NPs are synthesized via the microwave (MW)-assisted polyol route[17, 6, 38, 10, 40] in a CEM Discover SP MW reactor.
 First, the amount of is weighed into the reaction vial.
@@
 TMAH is consistently used at a molar ratio of relative to the Cu precursor.
-The reaction vessel is then placed in the microwave reactor, with the head space filled with inert argon up to a pressure of 5 bar, and is first stirred for minutes at to ensure complete dissolution of the precursor.
+The reaction vessel is then placed in the microwave reactor, with the head space filled with inert argon up to a pressure of 5 bar, and is first stirred for minutes at C to ensure complete dissolution of the precursor.
 Afterwards, the temperature is raised to the desired reaction temperature with continuous stirring and maintained for the desired reaction time.
 After the reaction time has passed, the reaction is quenched by actively cooling the vessel with compressed air.
-A brown-red suspension is obtained, which is typical for Cu NPs.[blosi:JournalNPResearch2010] To collect the Cu NPs, the reaction mixture is diluted with ml of ethanol and centrifuged at rpm for minutes.
+A brown-red suspension is obtained, which is typical for Cu NPs.[6] To collect the Cu NPs, the reaction mixture is diluted with ml of ethanol and centrifuged at rpm for minutes.
 Subsequently, the supernatant is removed, and the resulting NPs are again dispersed in ml of ethanol.
@@
 2.2 Computational modeling
-Two modeling tasks are performed on the data set: a regression task based on ensemble regression ML, implemented in the AMADEUS framework,[AMADEUS2020, vanpoucke:JournalApplPhys2020] and these results are compared with classical DoE[shahriari:ProceedingsIEEE2015], to gain insight in the potential of ML.
+Two modeling tasks are performed on the data set: a regression task based on ensemble regression ML, implemented in the AMADEUS framework,[43, 29] and these results are compared with classical DoE[36], to gain insight in the potential of ML.
 2.2.1 Ensemble regression models
-Since ML using small data sets is very sensitive to the selected training data, and hyperparameter tuning suffers from this volatility[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022], we make use of the scikit-learn-based AMADEUS framework which was designed to deal with small data sets (Figure 2).[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022, pedregosafabianJournalMLResearch2011, AMADEUS2020] Pasting-type ensemble models consisting of base instances are created, using either linear or polynomial base models in combination with LASSO regularization.[tibshirani:RoyalStatSocB1996] Each of the instances is trained on a different randomly selected 80/20 split (in-bag/out-of-bag (OOB)) of the complete data set.
+Since ML using small data sets is very sensitive to the selected training data, and hyperparameter tuning suffers from this volatility[29, 42], we make use of the scikit-learn-based AMADEUS framework which was designed to deal with small data sets (Figure 2).[29, 42, 16, 43] Pasting-type ensemble models consisting of base instances are created, using either linear or polynomial base models in combination with LASSO regularization.[39] Each of the instances is trained on a different randomly selected 80/20 split (in-bag/out-of-bag (OOB)) of the complete data set.
 The resulting final ensemble model is a single base instance equal to the average of the 100 base model instances.
@@
 1.
-Data set A: All data points are included in the data set, and the predicted particle size is the smallest particle size present in the DLS particle-size distribution.
+Data set A: All data points are included in the data set, and the predicted particle size is the smallest particle size present in the DLS particle-size distribution. (cf. SI Table VI)
 2.
 Data set B: All data points are included in the data set, and the predicted particle size corresponds to the peak with the highest percentage contribution in the DLS particle-size distribution.
+(cf.
+SI Table VII)
 3.
-Data set C: Only the data points with a mono-modal DLS particle-size distribution are included in the data set.
+Data set C: Only the data points with a mono-modal DLS particle-size distribution are included in the data set. (cf. SI Table VIII)
 For each of these data sets, ensembles of polynomial regression models of order to with LASSO regularization () are trained.
@@
 Therefore, no features above degree three are considered, further reducing overfitting.
-A further measure employed to reduce overfitting, is consideration of what one could call ensemble importance.[vanpoucke:PolymerInt2022, VerdingP:AdvMaterTech2025] The AMADEUS framework tracks how often a polynomial feature is retained (i.e.
+A further measure employed to reduce overfitting, is consideration of what one could call ensemble importance.[42, 45] The AMADEUS framework tracks how often a polynomial feature is retained (i.e.
 has non-zero coefficient) within the ensemble, after regularization of the features in each of the model instances.
@@
 2.2.2 Classical statistical methods (DoE)
-Using DoE[shahriari:ProceedingsIEEE2015] software JMP Pro (SAS Institute Inc.), a linear regression model is fitted for data set C, including terms up to the second degree.
+Using DoE[36] software JMP Pro (SAS Institute Inc.), a linear regression model is fitted for data set C, including terms up to the second degree.
 Insignificant terms are removed until only significant terms remain.
@@
 The performance of the model is determined by calculating the expected size for all data points using the equation obtained, and thus should be compared to quality estimates on the in-bag data set.
+2.2.3 Classification analysis
+A binary classification is performed on the particle size using data set A.
+(cf.
+SI Table VI) The aim is to predict whether the resulting particles will be large or small based on the synthesis parameters.
+Particles are defined as large if they are larger than the median size in the data set; if they are smaller than the median size, they are defined as small.
+For this classification task, the performance of complex LLMs is compared with that of a simple random forest classifier using accuracy and kappa () as performance metrics.
+indicates the agreement between predictions and actual labels, but corrects for coincidental agreements.
+It therefore indicates how well the model performs compared to random guessing.
+If there is a perfect match between predictions and actual labels, while indicates a performance equal to random guesses.
+If the model performs even worse than random guessing.
+Two different LLMs are trained and compared with each other for the binary classification task: GPT-J and LLAMA 3.1.[25, 41], to predict the outcome as large particles (‘1’) or small particles (‘0’) based on the experimental input parameters.
+The LLMs are fine-tuned using the following prompt as input:
+‘What is the size of nanoparticles synthesized with a mol/litre Copper Acetate precursor concentration, a mol/liter tetramethylammonium hydroxide concentration, heated for minutes at degrees celcius?’
+In this prompt, acts as a placeholder for the reaction parameter value.
+The LLM predicts ‘0’ or ‘1’ corresponding to experimentally small or large NPs, respectively.
+The training data set is limited to a maximum of examples, while the test data set remained constant across all runs (e.g., examples).
+The number of epochs[12] is first evaluated at , , , and .
+For each combination of training size and epoch count, five independent runs are performed to compute average performance metrics.
+GPT-J is used for this screening.
+Because the task involves binary classification of a balanced data set, any accuracy above % indicates performance better than random guessing.
+In comparison to the LLMs, an ensemble approach (e.g., random forest models) is used for binary classification.
+Automatic hyperparameter selection is used to determine the optimal value for the number of trees within a forest (), and the maximum depth of a tree ().
+Based on this hyperparameter tuning, is set to and to 3.
+With these hyperparameters, a single random forest is trained on a random subset of the data (80/20 train/test split).
+Repeating this process ten times, shows a high dependency on which specific data points are include in the training (and test) data, in line with previous observations with the volatility of results for small data ML.[29] Our results show the models accuracy varying between and .
+Because of this, an ensemble of random forests is trained, each random forest on a new random subset of the data.
+Each generated model is tested on its OOB data, and these predictions are logged.
+After training and evaluating all models, the OOB predictions of all models are used to obtain a final classification through hard voting.
+For the OOB predictions, the number of times they are classified as small (‘0’) and the number of times as large (‘1’) are counted for each data point.
+The class with the majority of votes will be the final predicted class.
 3 Results and Discussion
@@
 The remaining samples show a mono-modal particle-size distribution.
-Analysis of the DLS intensity-size plots allows us to identify the smallest particle size present.
+(cf.
+SI Figures 9, 10 and 11) Analysis of the DLS intensity-size plots allows us to identify the smallest particle size present.
 The median of these particle sizes is nm over all samples.
+(cf.
+SI Figure 7)
 Based on the SPR location, the effective formation of Cu metal () NPs is verified for the different samples.
-The different UV-Vis spectra show a clear SPR peak between nm and nm for most samples, which is a strong indication of the presence of NPs.[blosi:JournalNPResearch2010] Samples , , and do not show a clear SPR feature, which may be due to low particle concentration, oxidation[vargas-urbano:Materials2022], or the presence of ultra small particles (), which do not exhibit a well-defined SPR; for the other samples, at least some is present.
-Given the limited yield of the syntheses, no X-ray diffraction measurements could be performed to provide a conclusive answer, but the combination of the red-brown color of the solution and the presence of a clear SPR peak in the UV-Vis spectra strongly indicate that NPs have indeed formed.
+The different UV-Vis spectra show a clear SPR peak between nm and nm for most samples, which is a strong indication of the presence of NPs.[6] Samples , , and do not show a clear SPR feature, which may be due to low particle concentration, oxidation[44], or the presence of ultra small particles (), which do not exhibit a well-defined SPR; for the other samples, at least some is present.
+Given the limited yield of the syntheses, no X-ray diffraction measurements could be performed to provide a conclusive answer, but the combination of the red-brown color of the solution and the presence of a clear SPR peak in the UV-Vis spectra (cf.
+SI Figures 13 and 14 and SI Table IX) strongly indicate that NPs have indeed formed.
 To validate this conclusion, Selected Area Electron Diffraction analysis is conducted on sample 22 using a transmission electron microscope operated at an accelerating voltage of 120 kV.
 The resulting diffraction rings are indexed to the face-centered cubic structure of copper, with prominent rings corresponding to the (111), (200), and (220) planes.
-The primary diffraction rings are assigned the following spacings: Å for the (111) plane, Å for the (200) plane, and Å for the (220) plane.
+(cf.
+SI Figure 8) The primary diffraction rings are assigned the following spacings: Å for the (111) plane, Å for the (200) plane, and Å for the (220) plane.
 These values agree with the standard values for FCC copper (JCPDS card no.
-04-0836)[icdd2001], confirming the formation of NPs.
+04-0836)[24], confirming the formation of NPs.
 3.2 Quantitative size prediction
@@
 As first generation ensemble model, we consider the complete third-degree polynomial model with LASSO regularization ().
-This model has an of , and RMSE and MAE of nm and nm, respectively.
-Despite the very good performance of this first generation model, its main disadvantage is the use of an extensive set of features.
+(cf.
+SI Figure 15) This model has an of , and RMSE and MAE of nm and nm, respectively.
+(cf., Table 2) Despite the very good performance of this first generation model, its main disadvantage is the use of an extensive set of features.
 With our data set consisting of only 18 samples, the risk of overfitting is significant.
@@
 The smaller difference between training and OOB error indicates that the model is overfitting less and generalizes better.
-In the third generation model, the feature set is expanded with four additional features, aimed at the investigation of the influence of non-polynomial features (based on domain knowledge) on the model performance.
+(cf.
+SI Figure 16)
+In the third generation model (cf.
+SI Figure 17), the feature set is expanded with four additional features, aimed at the investigation of the influence of non-polynomial features (based on domain knowledge) on the model performance.
 1.
@@
 This more complex feature set, gives rise to a slight decline in the performance of the model due to stronger overfitting, leading to poorer generalization.
-Given the small difference in performance between this and the previous generation, a comprehensive analysis of the ensemble and dominance importance[Azen:PsyMeth2003, Ceder:ChemMater2022] of the features is carried out.
+(cf., Table 2) Given the small difference in performance between this and the previous generation, a comprehensive analysis of the ensemble and dominance importance[3, 22] of the features is carried out.
 Based on this analysis features are selected, aimed at a good balance between sufficient complexity for good predictive quality and a reduced size while retaining good generalization, which resulted in our fourth generation model.
-The dominance importance and ensemble importance are considered in an iterative fashion, as done in previous work of one of the authors.[vanpoucke:PolymerInt2022, VerdingP:AdvMaterTech2025] An exhaustive set of sub-models are trained, with one, two or three features (from the full feature list of the third generation model) systematically disabled.
+The dominance importance and ensemble importance are considered in an iterative fashion, as done in previous work of one of the authors.[42, 45] An exhaustive set of sub-models are trained, with one, two or three features (from the full feature list of the third generation model) systematically disabled.
 This allows us to determine the influence of each individual feature on the model’s performance.
@@
 Therefore, only the six features with an average ensemble importance % are kept for the fourth generation (cf.
-SI Table IX).
+SI Table X).
 | 1. T 2. 3. | 4. 5. 6. |
@@
 Because the size of the model has been greatly reduced, the MAE for the OOB predictions decreases from nm for the third generation to nm for the fourth generation, a marked improvement.
-Since the average of the ensemble model is equivalent to a single model instance with averaged coefficients [vanpoucke:JournalApplPhys2020], our fourth generation model can be formulated as follows:
+Since the average of the ensemble model is equivalent to a single model instance with averaged coefficients [29], our fourth generation model can be formulated as follows:
 In Equation 1, all feature values are standardized prior to training, meaning the coefficients apply to standardized inputs (cf.
-SI Table X).
+SI Table XI).
 When standardized experimental feature values are inserted into Equation 1, the model directly returns the predicted particle size in nm.
@@
 Because individual data points have a large impact in small datasets, models trained on the full dataset are highly sensitive to noise and outliers.
-Training multiple model instances on random subsets and averaging their predictions reduces this sensitivity, yielding more accurate and robust models.[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022] On the other hand, the classical regression model consists of only two terms containing the precursor concentration, compared to the six features in the ensemble regression model, which results in a simpler model and therefore lower computational cost.
+Training multiple model instances on random subsets and averaging their predictions reduces this sensitivity, yielding more accurate and robust models.[29, 42] On the other hand, the classical regression model consists of only two terms containing the precursor concentration, compared to the six features in the ensemble regression model, which results in a simpler model and therefore lower computational cost.
 Comparison of the ML and DoE models indicates that Cu concentration is an important factor, as both contain it as a polynomial term.
@@
 Since the DoE model only retained two features, we attempted to further reduce the fourth generation model by removal of additional features, based on dominance importance analyses (cf.
-SI Table XI).
+SI Table XII).
 First was removed, resulting in a slight performance reduction.
@@
 For three different target sizes ( nm, nm and nm) Equation 2 is used to suggest two possible syntheses routes for each (cf.
-SI Table XII).
+SI Table XIII).
 The suggested syntheses are performed using the same experimental procedure used for creating the original data set samples.
@@
 This shows that the DoE model exhibits a flatter, less pronounced curvature in the explored parameter space compared to the ML model.
-4 Conclusion
+3.3 Classification models
+In addition to the regression model, also a binary classification model is created.
+The goal of this model is to determine whether the obtained Cu NPs will be large (‘1’) or small (‘0’), based on the synthesis parameters.
+We compared the performance of complex LLMs, trained using transfer learning to a classic random forest model, trained from scratch.
+Early experiments with epochs are unsuccessful.
+The models failed to learn the prompt/completion structure, producing invalid outputs for the test data (e.g., numerical artifacts such as ) instead of valid binary labels (‘0’ or ‘1’), preventing the calculation of meaningful performance metrics.
+With a higher number of epochs, however, the fine-tuned model consistently generated valid predictions (‘0’ or ‘1’).
+(cf.
+SI Figure 12)
+For LLM-based classification, there is a clear difference between the performance of GPT-J and LLAMA.
+Where GPT-J performs slightly worse than random guessing, with an accuracy of and a (Table 4), the use of LLAMA 3.1 results in an increase in accuracy to and (Table 4).
+Although this is a significant performance improvement, the overall performance is still relatively low.
+One of the possible reasons for this limited performance is the very limited size of the data set.
+Small data sets already pose major challenges for classic ML, such as regression and decision trees, but the use of such small data sets is particularly challenging for complex models as LLMs.
+The classic random forest classifier has an overall performance between the two LLMs, with an accuracy of and a (Table 4).
+In this case, the performance is therefore slightly better than random guessing.
+The relatively limited difference in performance between LLMs and classic random forests raises questions about the usefulness of complex neural network based architectures in the context of limited lab-scale studies.
+It is clear that these require a much larger data set to achieve proper performance.[41] When the data set is limited, as in the current work, there is hardly any difference in their performance compared to random forests.
+Considering that LLMs are much more complex than classical models makes them less efficient with regard to the computational cost and thus energy requirements.
+| Accuracy | ||
+| GPT-J | ||
+| LLAMA 3.1 | ||
+| Random forest |
+4 Conclusions
 Despite the small size of the data set, it is possible to predict the particle size of Cu NPs quantitatively based on synthesis parameters using classical ensemble regression ML, achieving an of and an MAE of nm.
@@
 For two of the three target sizes, the suggested syntheses closely matched the experimental results, demonstrating that ML can effectively guide experimental synthesis even with a minimal data set.
+We additionally explored the applicability of LLMs for synthesis prediction in this small-data regime.
+While LLMs offer conceptual advantages, such as operating directly on recipe-style inputs without explicit feature engineering, our results indicate that they fail to identify reliable relationships between synthesis parameters and synthesis outcome from limited experimental data.
+While larger data sets might improve LLM performance, generating such data would conflict with the objective of data-efficient synthesis optimization in this current lab-scale experimental setting.
+These findings indicate that, for lab-scale NP synthesis with data sets typical in experimental materials research, carefully curated small data combined with interpretable classical ML models provides a more reliable approach than deploying more complex LLMs.
 Although this work focused on Cu NPs as a model system, the approach is broadly applicable to other materials.
@@
 Overall, this strategy contributes to more efficient, interpretable, and sustainable materials development.
-Acknowledgment
+Author contributions
+Brent Motmans: Data curation, Formal analysis, Investigation, Software, Writing–original draft, Writing–review & editing.
+Digvijay Ghogare: Supervision, Writing–review & editing.
+Thijs G.
+I.
+van Wijk: Supervision, Writing–review & editing.
+Joren Van Herck: Formal analysis, Software, Writing–review & editing.
+Pieter De Meyer: Formal analysis, Software, Writing–review & editing.
+Berend Smit: Supervision, Writing–review & editing.
+An Hardy: Conceptualization, Supervision, Writing–review & editing.
+Danny E.
+P.
+Vanpoucke: Conceptualization, Supervision, Writing–review & editing.
+Conflicts of interest
+There are no conflicts to declare.
+Data availability
+All data supporting this article are included in the manuscript and Supplementary Information.
+Acknowledgements
+The authors acknowledge Sander Stulens for performing the selected area electron diffraction (SAED) measurements.
+This work was financially supported by the Special Research Fund (BOF) via the BOF project, Belgium R15949.
 The computational resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government–department WEWIS.